Find the slope of the line perpendicular to the line joining point1(8,1) and point 2(2,4)?
Answer by mananth(16946)

x1 y1 x2 y2
8 * 1 2 4

slope m = (y2-y1)/(x2-x1)
( 4 - 1 )/( 2 - 8 )
( 3 / -6 )
m= - 1/ 2
The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be 2